# An Agentic Operationalization of DISARM for FIMI Investigation on Social Media

### v1

@@ Abstract
The interoperability of data and intelligence across allied partners and their respective end-user groups is considered a foundational enabler to the collective defense capability–both conventional and hybrid–of NATO countries. Foreign Information Manipulation and Interference (FIMI) and related hybrid activities are conducted across various societal dimensions and infospheres, posing an ever greater challenge to the characterization of threats, sustaining situational awareness, and response coordination. Recent advances in AI have further led to the decreasing cost of AI-augmented trolling and interference activities, such as through the generation and amplification of manipulative content. Despite the introduction of the DISARM framework as a standardized metadata and analytical framework for FIMI, operationalizing it at the scale of social media remains a challenge. We propose a framework-agnostic agent-based operationalization of DISARM to investigate FIMI on social media. We develop a multi-agent pipeline in which specialized agentic AI components collaboratively (1) detect candidate manipulative behaviors, and (2) map these behaviors onto standard DISARM taxonomies in a transparent manner. We evaluated the approach on two real-world datasets annotated by domain practitioners. We demonstrate that our approach is effective in scaling the predominantly manual and heavily interpretive work of FIMI analysis, providing a direct contribution to enhancing the situational awareness and data interoperability in the context of operating in media and information-rich settings.
@@ I Introduction
Foreign Information Manipulation and Interference (FIMI) is the term adopted by the EU’s External Action Service to describe state-sponsored disinformation and influence campaigns. Unlike generic “fake news,” FIMI is defined by coordinated malign behaviors often orchestrated by foreign actors with the intention of harming or destabilizing societies. Hostile actors are increasingly using Artificial Intelligence (AI) to supercharge their influence efforts. In addition to bots that amplify messages with inhuman speed, generative AI provides a low-cost option [14] to create inauthentic content and increase the scale of FIMI activities [3].
The Disinformation Analysis and Risk Management (DISARM) framework is an open-source taxonomy designed to standardize the identification and analysis of influence operations. The framework organizes the behaviors of threat actors into distinct Tactics, Techniques, and Procedures (TTPs), creating a common vocabulary for researchers, policymakers, and defenders. By systematically mapping the operational steps of a campaign, it facilitates cross-sector data sharing, enables more precise attribution of adversarial behaviors, and supports the development of evidence-based effective countermeasures against FIMI [19].
Despite the analytical value of DISARM, operationalizing it at the social-media scale remains difficult. Mapping TTPs typically requires labor-intensive and expert-driven coding of large, multilingual, and rapidly evolving datasets. This creates a gap between the conceptual promise of the framework and its day-to-day usability for timely detection, attribution, and response.
Agentic AI constitutes a leap in artificial intelligence systems, evolving beyond static rules-based tools into autonomous systems, suitable for pursuing complex, open-ended goals with minimal human intervention. Characterized by advanced reasoning, adaptive decision-making, and the capacity to decompose high-level objectives into executable sub-tasks, Agentic AI can operate effectively within dynamic and unstructured environments [1].
In this work, we propose an agent-based operationalization of DISARM to investigate FIMI on social media. We develop a multi-agent pipeline in which specialized agentic AI components collaboratively (1) detect candidate manipulative behaviors, and (2) map these behaviors onto DISARM TTPs in a transparent manner. We evaluate the approach on two real-world datasets annotated by FIMI domain practitioners. This allows us to test whether autonomous agents can recover known TTP patterns, how robustly they generalize across platforms and topics, and where human oversight remains essential.
We make three contributions to the practice of FIMI investigation:
@@
    Technique-guided agent architecture: We present a modular agent architecture that operationalizes DISARM as an executable investigative workflow, featuring technique-guided hypothesis generation where every discovered pattern maps to a specific TTP, and subprocess relay enabling consistent execution across 15 investigation rounds.
@@
    Verifiable evidence framework: We introduce atomic evidence decomposition–breaking complex findings into discrete, verifiable claims–combined with statistical validation, providing reproducible pass/fail criteria for LLM-generated evidence.
@@ II Rationale to military application
For military and defense organizations, the information environment is an operational space within the Multi-Domain Operations (MDO) paradigm, closely intertwined with, but not reducible to, the cyber domain. FIMI is increasingly treated as a key instrument of hybrid warfare, where information manipulation is used alongside other instruments of power to achieve strategic effects [18]. Because such campaigns can generate strategic effects below the threshold of armed conflict and at high tempo, defensive FIMI investigation must be scalable, structured, and aligned with alliance practices in order to support timely and proportionate responses. From a doctrinal perspective, FIMI defense maps naturally to the classical OODA loop (Observe-Orient-Decide-Act). Rapid observation of emerging narratives, robust orientation through structured interpretation, timely decision on proportional responses, and coordinated action across stakeholders are prerequisites to maintain initiative in the cognitive domain. Agentic AI provides a mechanism to compress the observe-orient phases by automating continuous analysis [^1] We do not focus on agentic collection in this work, while DISARM supplies a stable semantic backbone to ensure that machine-produced outputs remain interpretable and operationally meaningful [10]. An agent-based operationalization of DISARM therefore offers a concrete pathway to integrate FIMI defense into Command, Control, Intelligence, Surveillance, and Reconnaissance (C2ISR) processes.
@@ II-A Situational Awareness
Military decision-making depends on shared situational awareness, increasingly understood as multi-domain awareness that includes the information environment [5, 11]. In FIMI contexts, awareness requires persistent monitoring of social media networks at scale, rapid detection of coordinated inauthentic behavior (CIB), and mapping of narrative spread across platforms and languages [6]. The speed and volume of AI-supported influence operations mean that delays in detection can directly translate into strategic disadvantage. An agentic DISARM supports situational awareness by allowing autonomous agents to collect signals, identify candidate manipulative behaviors, and continuously map these behaviors to DISARM TTPs. This produces structured, time-sensitive indicators of adversarial activity rather than unorganized streams of alerts. In practice, it allows human analysts and commanders to move from ad hoc monitoring to an always-on early warning posture, improving both the timeliness and the granularity of FIMI threat pictures. Importantly, this awareness is actionable because it is expressed in a standardized behavioral taxonomy, facilitating prioritization and response planning.
@@ II-B Human-Machine Teaming
A recent CSIS study [4] on agentic warfare argues that Napoleonic‑style, centralized staffs are too slow and brittle for conflicts where AI agents compress decision cycles from days to minutes and adversaries deliberately target command networks—thereby necessitating smaller, adaptive, AI‑enabled staffs that preserve human judgment while operating at machine speed. In this context, our proposed workflow treats agentic systems as accelerators of staff work rather than replacements for FIMI analysts and responders: agents continuously monitor, decompose tasks, surface hypotheses and organize evidence, while expert analysts validate, contextualize and interpret [13]—especially in inherently interpretive FIMI workflows that require assessing intent, cultural cues and proportionality where full automation is neither desirable nor feasible. Operationalizing DISARM with agents provides the doctrinal scaffold analogous to CSIS recommendations for resilient human–machine teaming: instead of opaque labels, agents must justify each assignment as an instance of specific TTPs within an explicit taxonomy, yielding explainable, traceable outputs from raw content to analytic conclusions, reducing automation bias, and enabling auditability. In short, autonomous agents extend analyst capacity at machine speed, while DISARM‑aligned outputs maintain transparency, accountable judgment, and doctrinal coherence—exactly the combination required for adaptive staff models [4] envisioned for future military operations.
@@ II-C Interoperability
In alignment with NATO’s strategic move toward a ”digitally enabled, multi-domain Alliance,” an agentic DISARM pipeline serves as a critical ”interoperability by design” solution that addresses the technical, procedural and human dimensions of collective defense [12]. By utilizing the DISARM framework as an ontological foundation, the pipeline establishes a procedural common language that standardizes investigative outputs across national contexts, effectively reducing the ”interpretive drift” that often hinders joint responses to foreign influence. This semantic standardization enables technical interoperability through the multi-layer fusion [16, 8] of incident data. From a human interoperability perspective, the efficiency gain through data fusion in turn facilitates the emergence of an agentic military staff [4] to process information at a speed and across scales that can exceed those of traditional staffing structures. Ultimately, by automating the cross-validation and attribution of transnational threats, the ”agent-agumented” pipeline can transform fragmented national-level analyses into a unified, shareable ecosystem, enabling the collective defense against transnational information threats.
@@ III Methods
This section presents our agent-based pipeline for operationalizing the DISARM framework developed for FIMI investigations. The pipeline, built on BMAD-METHOD (Breakthrough Method for Agile AI-Driven Development)[9], enables systematic discovery and statistical verification of IO indicators through iterative hypothesis generation and testing.
@@ III-A Pipeline overview
Our approach addresses three key challenges in automating FIMI detection: (1) mapping discovered anomalies to standardized IO taxonomies, (2) ensuring statistical rigor in evidence validation, and (3) maintaining full traceability from raw data to verified findings. The pipeline operates in two phases as illustrated in the upper and lower parts of Figure 1:
@@
  • Deferred anomaly detection: Initial exploration provides data overview without anomaly detection, deferring specific findings to technique-guided iterations. This ensures every discovered pattern maps to a specific DISARM technique.
  • Full history feedback: Each iteration accesses the complete history of prior findings, enabling informed technique selection and pattern recognition across iterations.
  • Evidence-based verification: All claims undergo statistical validation using ground truth labels or annotations, with explicit pass/fail criteria based on effect size thresholds.
  • Natural Language specification: Workflows, tasks, and agent behaviors are defined using human language rather than traditional code, enabling rapid adaptation to diverse OSINT scenarios without software development expertise.
@@ III-B1 Chinese campaign on X focused on the coordinated attacks on Guo Wengui in 2017/2019
We selected the China 1 and China 2 datasets introduced by Seckin [17] as part of a large collection of labeled information operations (IO) datasets. These datasets originate from transparency disclosures on a social media platform and correspond to two distinct but thematically related IO campaigns attributed to Chinese state-backed actors.
In the original datasets, the platform explicitly identified and verified that the IO accounts participate in coordinated information operations. To enable supervised evaluation of IO detection methods, the authors complement the IO data with a control dataset representing organic user activity. Control accounts are defined as accounts not identified as IO by the platform but that discuss similar topics during the same time periods as the IO campaigns, ensuring topical and temporal overlap between IO and non-IO activity. Control accounts are collected using a hashtag-based matching strategy: hashtags used by IO accounts on specific dates are used to identify other accounts posting on the same dates with at least one of those hashtags.
In this work, we merge China 1 and China 2 into a single dataset. We then apply additional processing steps to obtain a simplified dataset focused on the coordinated attacks on Guo Wengui, a Chinese businessman and political dissident whose activities and public statements have made him a recurrent target of state-aligned online campaigns. First, we restrict the data to posts in Mandarin Chinese related to the campaign using a curated list of keywords and hashtags associated with narratives and attacks targeting Guo Wang. This focuses on the central content to the information operation while preserving both IO and control users. Second, we apply a uniform text-cleaning pipeline to both IO and control posts, including the removal of URLs, mentions, hashtags, emojis, and extraneous whitespace, while preserving Chinese characters. The posts are tokenized, and only posts with at least 10 valid tokens are retained to eliminate short or noisy messages.
The resulting dataset consists of 617 IO accounts and 157 organic accounts, comprising 40,863 clean Chinese-language posts (38,927 IO posts and 1,936 organic posts) drawn from both platform-verified IO accounts and organic control accounts, sharing a common topical and temporal context.
@@ III-B2 Russian campaign on Telegram for 2025 Moldova Election
This dataset captures coordinated and automated political commenting activity on Moldovan Telegram channels during a period of heightened electoral interference. The data were collected and verified by the OpenMinds research team as part of its investigation into Russian information operations targeting Moldova (see published report at [15]).
@@
Bot identification focused exclusively on user-generated comments posted in these open comment sections between July 1 and September 15, 2025. Within this period, 462 unique accounts that exhibited bot-like behavior were identified. These accounts collectively produced 62,229 comments on the selected channels. All identified accounts were manually verified and classified as bots after a structured review process.
Bot detection relied on a combined methodology that integrated algorithmic indicators and expert human judgment. Algorithmic signals included unusually high commenting volume, temporal regularity, repetitive posting patterns, and coordinated activity across multiple channels. Subsequently, these signals were assessed by manual inspection to exclude high-activity organic users and ensure that only coordinated or automated behavior was retained in the final data set.
Following the publication of the initial study, data collection was extended until the end of September 2025 to capture activity on election day. The final dataset comprises 595,814 total comments, produced by 30,297 unique accounts, of which 584 accounts were verified as bots. Bot accounts generated 86,448 comments, representing 14,5% of all comments in the data set. The final released dataset therefore represents a high-confidence corpus of bot-driven political commenting activity on Moldovan Telegram, grounded in manual verification and consistent behavioral classification.

### v2

@@ Abstract
The interoperability of data and intelligence across allied partners and their respective end-user groups is considered a foundational enabler of the collective defense capability—both conventional and hybrid—of NATO countries. Foreign Information Manipulation and Interference (FIMI) and related hybrid activities are conducted across various societal dimensions and infospheres, posing an ever greater challenge to threat characterization, sustained situational awareness, and response coordination. Recent advances in AI have further reduced the cost of AI-augmented trolling and interference activities, such as through the generation and amplification of manipulative content. Despite the introduction of the DISARM framework as a standardized metadata and analytical framework for FIMI, operationalizing it at the scale of social media remains a challenge. We propose a framework-agnostic, agent-based operationalization of DISARM to investigate FIMI on social media. We develop an agent coordination pipeline in which specialized agentic AI components collaboratively (1) detect candidate manipulative behaviors and (2) map these behaviors onto standard DISARM taxonomies in a transparent manner. We evaluate the approach on two real-world datasets annotated by domain practitioners. Our results show that the approach is effective in scaling the predominantly manual and heavily interpretive work of FIMI analysis—including uncovering more than 30 previously undetected Russian bot accounts during manual analysis—and provides a direct contribution to enhancing situational awareness and data interoperability in the context of operating in media- and information-rich settings.
@@ I Introduction
Foreign Information Manipulation and Interference (FIMI) is the term adopted by the EU’s External Action Service to describe state-sponsored disinformation and influence campaigns. Unlike generic “fake news,” FIMI is characterized by coordinated malign behaviors, often orchestrated by foreign actors with the intention of harming or destabilizing societies. Hostile actors are increasingly using artificial intelligence (AI) to supercharge their influence efforts. In addition to bots that amplify messages at inhuman speeds, generative AI provides a low-cost option[14] for creating inauthentic content and increasing the scale of FIMI activities[3].
The Disinformation Analysis and Risk Management (DISARM) framework is an open-source taxonomy designed to standardize the identification and analysis of influence operations. The framework organizes threat actor behaviors into distinct tactics, techniques, and procedures (TTPs), creating a common vocabulary for researchers, policymakers, and defenders. By systematically mapping the operational steps of a campaign, it facilitates cross-sector data sharing, enables more precise attribution of adversarial behaviors, and supports the development of effective, evidence-based countermeasures against FIMI [19].
Despite the analytical value of DISARM, operationalizing it at social media scale remains difficult. Mapping TTPs typically requires labor-intensive, expert-driven coding of large, multilingual, and rapidly evolving datasets. This creates a gap between the conceptual promise of the framework and its day-to-day usability for timely detection, attribution, and response.
Agentic AI represents a leap in artificial intelligence, evolving beyond static, rule-based tools into autonomous systems capable of pursuing complex, open-ended goals with minimal human intervention. Characterized by advanced reasoning, adaptive decision-making, and the ability to decompose high-level objectives into executable sub-tasks, Agentic AI can operate effectively in dynamic and unstructured environments [1].
In this work, we propose an agent-based operationalization of DISARM to investigate FIMI on social media. We develop an agent coordination pipeline in which specialized agentic AI components collaboratively (1) detect candidate manipulative behaviors and (2) map these behaviors onto DISARM TTPs in a transparent manner.
We evaluate the approach on two real-world datasets annotated by FIMI domain practitioners. This allows us to test whether autonomous agents can recover known TTP patterns, how robustly they generalize across platforms and topics, and where human oversight remains essential. We make three contributions to the practice of FIMI investigation:
@@
    Technique-guided agent architecture: We present a modular agent architecture that operationalizes DISARM as an executable investigative workflow, featuring technique-guided hypothesis generation, in which every discovered pattern maps to a specific TTP, and subprocess relay, which enables consistent execution across 15 investigation rounds.
@@
    Verifiable evidence pipeline: We introduce atomic evidence decomposition–breaking complex findings into discrete, verifiable claims–combined with statistical validation, providing reproducible pass/fail criteria for LLM-generated evidence.
@@ II Rationale to military application
For military and defense organizations, the information environment is an operational space within the Multi-Domain Operations (MDO) paradigm, closely intertwined with, but not reducible to, the cyber domain. FIMI is increasingly treated as a key instrument of hybrid warfare, in which information manipulation is used alongside other instruments of power to achieve strategic effects [18]. Because such campaigns can generate strategic effects below the threshold of armed conflict and at high tempo, defensive FIMI investigation must be scalable, structured, and aligned with alliance practices in order to support timely and proportionate responses. From a doctrinal perspective, FIMI defense maps naturally onto the classical OODA loop (Observe–Orient–Decide–Act). Rapid observation of emerging narratives, robust orientation through structured interpretation, timely decision on proportional responses, and coordinated action across stakeholders are prerequisites for maintaining initiative in the cognitive domain. Agentic AI provides a mechanism to compress the observe–orient phases by automating continuous analysis [^1] We do not focus on agentic collection in this work, while DISARM supplies a stable semantic backbone to ensure that machine-produced outputs remain interpretable and operationally meaningful [10]. An agent-based operationalization of DISARM therefore offers a concrete pathway to integrate FIMI defense into Command, Control, Intelligence, Surveillance, and Reconnaissance (C2ISR) processes.
@@ II-A Situational Awareness
Military decision-making depends on shared situational awareness, increasingly understood as multi-domain awareness that includes the information environment [5, 11]. In FIMI contexts, awareness requires persistent monitoring of social media networks at scale, rapid detection of coordinated inauthentic behavior (CIB), and mapping of narrative spread across platforms and languages [6]. The speed and volume of AI-supported influence operations mean that delays in detection can directly translate into strategic disadvantage. An agentic DISARM supports situational awareness by allowing autonomous agents to collect signals, identify candidate manipulative behaviors, and continuously map these behaviors to DISARM TTPs. This produces structured, time-sensitive indicators of adversarial activity rather than unorganized streams of alerts. In practice, it allows human analysts and commanders to move from ad hoc monitoring to an always-on early-warning posture, improving both the timeliness and the granularity of FIMI threat pictures. Importantly, this awareness is actionable because it is expressed in a standardized behavioral taxonomy, facilitating prioritization and response planning.
@@ II-B Human-Machine Teaming
A recent CSIS study [4] on agentic warfare argues that Napoleonic‑style, centralized staffs are too slow and brittle for conflicts in which AI agents compress decision cycles from days to minutes and adversaries deliberately target command networks—thereby necessitating smaller, adaptive, AI‑enabled staffs that preserve human judgment while operating at machine speed. In this context, our proposed workflow treats agentic systems as accelerators of staff work rather than replacements for FIMI analysts and responders: agents continuously monitor, decompose tasks, surface hypotheses, and organize evidence, while expert analysts validate, contextualize, and interpret [13]—especially in inherently interpretive FIMI workflows that require assessing intent, cultural cues, and proportionality, where full automation is neither desirable nor feasible. Operationalizing DISARM with agents provides the doctrinal scaffold analogous to CSIS recommendations for resilient human–machine teaming: instead of opaque labels, agents must justify each assignment as an instance of specific TTPs within an explicit taxonomy, yielding explainable, traceable outputs from raw content to analytic conclusions, reducing automation bias, and enabling auditability. In short, autonomous agents extend analyst capacity at machine speed, while DISARM‑aligned outputs maintain transparency, accountable judgment, and doctrinal coherence—exactly the combination required for the adaptive staff models [4] envisioned for future military operations.
@@ II-C Interoperability
In alignment with NATO’s strategic move toward a ”digitally enabled, multi-domain Alliance,” an agentic DISARM pipeline serves as a critical ”interoperability-by-design” solution that addresses the technical, procedural, and human dimensions of collective defense [12]. By utilizing the DISARM framework as an ontological foundation, the pipeline establishes a procedural common language that standardizes investigative outputs across national contexts, effectively reducing the ”interpretive drift” that often hinders joint responses to foreign influence. This semantic standardization enables technical interoperability through the multi-layer fusion [16, 8] of incident data. From a human interoperability perspective, the efficiency gain through data fusion in turn facilitates the emergence of an agentic military staff [4] able to process information at a speed and across scales that can exceed those of traditional staffing structures. Ultimately, by automating the cross-validation and attribution of transnational threats, the ”agent-augmented” pipeline can transform fragmented national-level analyses into a unified, shareable ecosystem, enabling collective defense against transnational information threats.
@@ III Methods
This section presents our agent-based pipeline for operationalizing the DISARM framework, developed for FIMI investigations. The pipeline, built on BMAD-METHOD (Breakthrough Method for Agile AI-Driven Development)[9], enables the systematic discovery and statistical verification of IO indicators through iterative hypothesis generation and testing.
@@ III-A Pipeline overview
Our approach addresses three key challenges in automating FIMI detection: (1) mapping discovered anomalies to standardized IO taxonomies, (2) ensuring statistical rigor in evidence validation, and (3) maintaining full traceability from raw data to verified findings. The pipeline operates in two phases, as illustrated in the upper and lower parts of Figure 1:
@@
  • Deferred anomaly detection: The initial exploration phase focuses on establishing a neutral understanding of the structure, scale, and behavioral baselines without attempting to identify anomalies or assign interpretive meaning. This separation prevents premature, framework-driven interpretation and allows DISARM mapping to emerge only during later, technique-guided iterations, when hypotheses are explicitly formulated and tested.
  • Full history feedback: Each iteration accesses the complete history of prior findings, enabling informed technique selection and contextual reasoning across iterations.
  • Evidence-based verification: All claims undergo statistical validation using ground-truth labels or annotations, with explicit pass/fail criteria based on effect-size thresholds.
  • Natural Language specification: Workflows, tasks, and agent behaviors are defined using natural language rather than traditional code, enabling rapid adaptation to diverse OSINT scenarios without requiring software development expertise.
@@ III-B1 Chinese campaign on X focused on the coordinated attacks on Guo Wengui in 2017/2019
We selected the China 1 and China 2 datasets introduced by Seckin [17] from a large collection of labeled information operations (IO) datasets. These datasets originate from transparency disclosures on a social media platform and correspond to two distinct but thematically related IO campaigns attributed to Chinese state-backed actors.
In the original datasets, the platform explicitly identified and verified that the IO accounts participated in coordinated information operations. To enable supervised evaluation of IO detection methods, the authors complemented the IO data with a control dataset representing organic user activity. Control accounts are defined as accounts not identified as IO by the platform but that discuss similar topics during the same time periods as the IO campaigns, ensuring topical and temporal overlap between IO and non-IO activity. Control accounts are collected using a hashtag-based matching strategy: hashtags used by IO accounts on specific dates are used to identify other accounts posting on the same dates with at least one of those hashtags.
In this work, we merge China 1 and China 2 into a single dataset. We then apply additional processing steps to obtain a simplified dataset focused on coordinated attacks on Guo Wengui, a Chinese businessman and political dissident whose activities and public statements have made him a recurrent target of state-aligned online campaigns. First, we restrict the data to posts in Mandarin Chinese related to the campaign using a curated list of keywords and hashtags associated with narratives and attacks targeting Guo Wengui. This focuses on the central content of the information operation while preserving both IO and control users. Second, we apply a uniform text-cleaning pipeline to both IO and control posts, including the removal of URLs, mentions, hashtags, emojis, and extraneous whitespace, while preserving Chinese characters. The posts are tokenized, and only posts with at least 10 valid tokens are retained to eliminate short or noisy messages.
The resulting dataset consists of 617 IO accounts and 157 organic accounts, comprising 40,863 cleaned Chinese-language posts (38,927 IO posts and 1,936 organic posts) drawn from both platform-verified IO accounts and organic control accounts, sharing a common topical and temporal context.
@@ III-B2 Russian campaign on Telegram for 2025 Moldova Election
This dataset captures coordinated and automated political commenting activity on Moldovan Telegram channels during a period of heightened electoral interference. The data were collected and verified by the OpenMinds research team as part of its investigation into Russian information operations targeting Moldova (see published report in [15]).
@@
Bot identification focused exclusively on user-generated comments posted in these open comment sections between July 1 and September 15, 2025. During this period, 462 unique accounts that exhibited bot-like behavior were identified. These accounts collectively produced 62,229 comments on the selected channels. All identified accounts were manually verified and classified as bots following a structured review process.
Bot detection relied on a combined methodology that integrated algorithmic indicators and expert human judgment. Algorithmic signals included unusually high commenting volume, temporal regularity, repetitive posting patterns, and coordinated activity across multiple channels. These signals were subsequently assessed through manual inspection to exclude high-activity organic users and ensure that only coordinated or automated behavior was retained in the final data set.
Following the publication of the initial study, data collection was extended until the end of September 2025 to capture activity on election day. The final dataset comprises 595,814 total comments, produced by 30,297 unique accounts, of which 584 accounts were verified as bots. Bot accounts generated 86,448 comments, representing 14.5% of all comments in the data set. The final released dataset therefore represents a high-confidence corpus of bot-driven political commenting activity on Moldovan Telegram, grounded in manual verification and consistent behavioral classification.